OpenStudy (anonymous):

Help please. write the equation of the line through each point use slope intercept form(1 -1) parallel to y=2/5x-3 please show me step by step because I am trying to understand how to solve these.

OpenStudy (tkhunny):

You must have the "Point Slope" form. Have you?

OpenStudy (anonymous):

I do not have that

OpenStudy (tkhunny):

You need a point, usually designated \((x_{0},y_{0})\) In your case, this is (1,-1) Thus, \(x_{0} = 1\;and\;y_{0} = -1\) Make sense, so far? We're just defining things.

OpenStudy (anonymous):

Yes this part makes sense

OpenStudy (tkhunny):

You need a slope, usually designated m. In your case, we have a hint. "parallel to y=2/5x-3" This is in Slope-Intercept form. The slope of this line is the slope of the line we seek. Thus m = 2/5 There are various other methods of telling us the slope indirectly. With this hint, we are expected to know that parallel lines have the same slope. If a line is parallel to y = (2/5)x-3, then the slope is 2/4. Now how are we doing? Still hanging in there?

OpenStudy (anonymous):

yes i am understanding it so far

OpenStudy (tkhunny):

Now, you need the form. The "Point-Slope" form of a line looks like this: \((y-y_{0}) = m(x-x_{0})\) You know already \(x_{0},\;y_{0},\;and\;m\). Substitute these values in the the Form and see where ti leads.